# On the contraction properties of Sinkhorn semigroups

### v2

@@ Abstract
We develop a novel semigroup stability analysis based on Lyapunov techniques and contraction coefficients to prove exponential convergence of Sinkhorn equations on weighted Banach spaces. This operator-theoretic framework yields exponential decays of Sinkhorn iterates towards Schrödinger bridges with respect to general classes of $\phi$-divergences and Kantorovich-type criteria, including the relative entropy, squared Hellinger integrals, $\alpha$-divergences as well as weighted total variation norms and Wasserstein distances. To the best of our knowledge, these contraction inequalities are the first results of this type in the literature on entropic transport and the Sinkhorn algorithm.
@@
We also provide Lyapunov contractions principles under minimal regularity conditions that allow to provide quantitative exponential stability estimates for a large class of Sinkhorn semigroups. We apply this novel framework in a variety of situations, ranging from polynomial growth potentials and heavy tailed marginals on general normed spaces to more sophisticated boundary state space models, including semi-circle transitions, Beta, Weibull, exponential marginals as well as semi-compact models. Last but not least, our approach also allows to consider statistical finite mixture of the above models, including kernel-type density estimators of complex data distributions arising in generative modeling.
@@
Keywords: Entropic Optimal Transport, Sinkhorn semigroups, Schrödinger bridges, contraction coefficients, weighted total variation norms, Lyapunov functions, Kantorovich and entropy criteria, Wasserstein semi-distances.
Mathematics Subject Classification: 37M25, 49Q22, 47H09, 60J20; secondary: 37M25, 60J05, 94A17.
1 Introduction
1.1 Entropic optimal transport
@@
Consider locally compact Polish spaces $(\mathbb{X},\lambda)$ and $(\mathbb{Y},\nu)$ endowed with some locally bounded positive measures $\lambda$ and $\nu$. Consider some probability measures of the form
$\lambda_{U}(dx):=e^{-U(x)}~\lambda(dx)\quad\mbox{\rm and}\quad\nu_{V}(dy):=e^{-V(y)}~\nu(dy)$ (1)
@@
for some normalized potential functions $U:x\in\mathbb{X}\mapsto U(x)\in\mathbb{R}$ and $V:y\in\mathbb{X}\mapsto V(y)\in\mathbb{R}$. Let $\Pi_{U,V}$ be the set of coupling probability measures $\cal P$} on the product space $(\mathbb{X}\times\mathbb{Y})$, with prescribed first and second coordinate marginals $(\lambda_{U},\nu_{V})$. Denote $\cal Q$} a Markov transition from $\mathbb{X}$ into $\mathbb{Y}$ of the form
${\cal Q}(x,dy)=e^{-W(x,y)}~\nu(dy)\quad\mbox{\rm and set}\quad{\cal R}(y,dx):=e^{-W(x,y)}~\lambda(dx).$ (2)
In the latter expression, the transition potential $W(x,y)$ stands for some real valued function on $(\mathbb{X}\times\mathbb{Y})$. Consider the reference probability measure constructed as
${\cal P}_{0}(d(x,y)):=(\lambda_{U}\times{\cal Q})(d(x,y)):=\lambda_{U}(dx)~{\cal Q}(x,dy).$ (3)
@@
The entropic transport problem (or Schrödinger bridge problem) with respect to (w.r.t.) the reference measure ${\cal P}_{0}$ is formulated as
@@
$\mathbb{P}=\operatorname*{arg\,min}_{{\cal P}\,\in\,\Pi_{U,V}}\mbox{\rm Ent}({\cal P}~|~{\cal P}_{0}),\quad\mbox{\rm with the relative entropy}\quad\text{Ent}({\cal P}~|~{\cal P}_{0}):=\int\log\frac{d{\cal P}}{d{\cal P}_{0}}~d{\cal P}.$ (4)
@@
It is implicitly assumed there exists some ${\cal P}\in\Pi_{U,V}$ such that $\mbox{\rm Ent}({\cal P}~|~{\cal P}_{0})<\infty$. This condition ensures the existence of a Schrödinger bridge distribution $\mathbb{P}$ solving (4) (cf. the seminal article by Csiszár [csiszar-2], as well as Section 6 in the Lecture Notes by Nutz [nutz], and also the survey article by Léonard [leonard] and references therein). Let us note that the product distribution
@@
trivially belongs to $\Pi_{U,V}$. In addition, choosing ${\cal P}=(\lambda_{U}\times{\cal K})\,\in\,\Pi_{U,V}$ for some Markov transition ${\cal K}(x,dy)$ we have
$\frac{d{\cal P}}{d{\cal P}_{0}}(x,y)=\frac{d(\lambda_{U}\otimes\nu_{V})}{d{\cal P}_{0}}(x,y)\times\frac{d{\cal P}}{d(\lambda_{U}\otimes\nu_{V})}(x,y)$
which yields
$\mbox{\rm Ent}({\cal P}~|~{\cal P}_{0})=\int~W(x,y)~{\cal P}(d(x,y))+\mbox{\rm Ent}({\cal P}~|~\lambda_{U}\otimes\nu_{V})-\int V(y)\nu_{V}(dy).$
@@
Choosing a potential transition of the form $W(x,y)={\cal W}(x,y)/t$, for some parameter $t>0$, problem (4) is equivalent to the optimal entropic transport problem
@@
$\mathbb{P}=\operatorname*{arg\,min}_{{\cal P}\,\in\,\Pi_{U,V}}\left(\int~{\cal W}(x,y)~{\cal P}(d(x,y))+t~\mbox{\rm Ent}({\cal P}~|~\lambda_{U}\otimes\nu_{V})\right).$ (5)
@@
In the optimal transport literature, function $\cal W$} is sometimes called the cost function and the parameter $t>0$ is interpreted as a regularization parameter. In this context, the non-regularized problem corresponding to $t=0$ coincides with the Kantorovich optimal transport problem with cost function $\cal W$}. For a more detailed discussion on this subject we refer to [dobrushin2006perturbation, rachev1985monge, villani2008optimal] and the numerous references therein.
1.2 Background
@@
The Schrödinger bridge problem (4) is to find a random evolution with minimal relative entropy w.r.t. some reference physical model and with prescribed distributions of the initial and final time random states. This convex minimization problem is a regularized version of the Monge-Kantorovich optimal transport with a relative entropy constraint, see  [leonard, nutz, peyre] for recent surveys with numerous references on this subject.
@@
Optimal transport methodologies have become state-of-the-art tools in a variety of application domains, including generative modeling and machine learning [arjovsky, doucet-bortoli, kolouri, peyre, rioux2023semi, rioux2024entropic], statistical barycenter problems [agueh, andersen, bigot, cramer-2, cuturi-w, loubes], economy [bojitov], computer vision [dominitz, solomon], control theory [chen, chen-2], and many others.
Under mild integrability conditions, the Schrödinger bridge problem has a unique solution which can be expressed as the fixed point of Sinkhorn equations also referred to as the iterative proportional fitting procedure. For a more thorough discussion on the origins of Sinkhorn iterations, we refer to the early and pioneering articles [deming, fortet, kruithof, sinkhorn-2, sinkhorn-3] as well as the more recent studies [chen, cuturi, leonard-2].
Sinkhorn bridges are based on nonlinear integral transformations over the whole state space (cf. Section 2) and exact finite-dimensional solutions cannot be computed for general models. However, for finite state spaces the Sinkhorn algorithm is a well known simple procedure that repeatedly applies column and row normalization on matrix spaces [deming, fortet, kruithof, sinkhorn-2, sinkhorn-3]. For Gaussian models, Sinkhorn equations can also be solved analytically using sequentially Bayes conjugacy principles [adm-24, cramer-2, cramer-3, janadi].
Most of the literature on the exponential convergence of Sinkhorn equations is concerned with finite state spaces [borwein, fienberg, sinkhorn-2, sinkhorn-3, soules] as well as compact state spaces or bounded cost functions using Hilbert projective metrics techniques [chen, deligiannidis, franklin, marino]. See [greco2023non] for a refined analysis of Sinkhorn iterates and their gradient on the $d$-dimensional torus as well as the more recent article [chizat2025sharper] on compact spaces and bounded cost functions.
As remarked in Section 3 of [cohen], the Hilbert metric dominates all $\phi$-divergences, thus we expect rather strong conditions to ensure the contraction of Sinkhorn transitions w.r.t. the Hilbert metric, see for instance [borwein, chen, deligiannidis, franklin] and Section 4.1 in the present article. Contraction estimates w.r.t. the Hilbert metric yield exponential decays w.r.t. $\phi$-divergences, including the total variation norm, the relative entropy and Wasserstein distances of any order. Nevertheless they can rarely be used to derive contraction estimates with respect to some $\phi$-divergence criteria.
At this point, we underline that even on totally bounded state spaces such as the open unit interval $\mathbb{X}=]0,1[=\mathbb{Y}$ the cost function $W$ as well as the potential functions $(U,V)$ may be unbounded and the reference Markov transition may have a unit Birkhoff contraction coefficient (i.e. equal to $1$), see for instance Section 3.2.1 and the series of illustrations discussed in Section 5.3 and Section 5.4.
The landmark example of quadratic costs on Euclidian spaces is another class of unbounded costs functions. In this context, the recent article [adm-24] also presents a refined stability analysis of Sinkhorn bridges for Gaussian models based on matrix Riccati difference equations. It is clearly out of the scope of this article to review all the contributions in this field and we simply refer to the recent book [peyre] as well as the more recent articles [adm-24, chiarini, durmus] and the references therein.
@@
Sub-linear convergence rates of Sinkhorn equations have also been developed in the articles [alts-2017, chak-2018, dvu-2017]. Linear rates on non-necessarily compact spaces were first obtained by Léger in [leger] using elegant gradient descent and Bregman divergence techniques (see also the recent articles [doucet-bortoli, karimi-2024]). Refined convergence rates at least one order faster for a general class of models including sub-Gaussian target marginals has also been developed in [promit-2022] under some exponential moments estimates on Schrödinger potentials.
There are very few and recent articles on the exponential convergence of Sinkhorn iterates for general models on non-compact spaces and unbounded cost functions that apply to Gaussian models, with the notable exception of three significant contributions [chiarini, durmus, eckstein2025hilbert].
The first exponential stability analysis of unbounded models and non necessarily compactly supported marginals is presented in the article [durmus]. In this study, the authors investigate quantitative contraction rates for target marginal distributions with an asymptotically positive log-concavity profile and quadratic symmetric cost functions associated with a sufficiently large regularization parameter. In our framework this particular class of cost functions corresponds to symmetric Gaussian reference transitions.
The approach taken in [durmus] relies on well known links between gradient and Hessian of Sinkhorn and Schrödinger potentials along Hamilton-Jacobi-Bellman equations with conditional expectations and covariances [chewi2023entropic, pooladian2021entropic], see also [del2025entropic]. The main conditions in [durmus] are expressed in terms of the integrated convexity profiles of marginal potentials. Extending ideas from [conforti2024weak] the proofs are based on the propagation of these convexity profiles based on diffusion reflection techniques introduced in [eberle2016reflection].
The exponential entropic decays presented in [durmus] were further developed and recently refined in [chiarini] to apply to all values of the regularization parameter and some classes of perturbation of log-concave models. The article [chiarini] has opened new perspectives in the convergence analysis for entropically regularized transport problems with quadratic symmetric costs. In the context of distance-like cost functions, one of the basic tool is a remarkable stability theorem, Theorem 1.1 in [chiarini], that allows to control the entropy of Schrödinger bridges by the entropy of their marginal distributions. However this stability approach requires to estimate the semi-concavity properties of generally unknown dual Schrödinger bridge transitions and it only applies to entropically regularized transport problems with symmetric quadratic-type costs. Thus, this framework excludes the important cases when the reference transition is associated with linear Gaussian transitions arising in Ornstein-Uhlenbeck-type diffusion generative models and denoising diffusions.
In this connection, the more recent article [del2025stability] complements the stability approach in  [chiarini] with a novel entropic continuity theorem that applies to more general classes of non necessarily symmetric transition potentials and log-concave type target marginals, including strongly convex at infinity potentials.
@@
The article [eckstein2025hilbert] is also based on a novel version of Hilbert’s projective metric for spaces of integrable functions of bounded growth. Extending results in [chen, deligiannidis] to unbounded cost functions, the recent article  [eckstein2025hilbert] presents exponential decays of these new classes of Hilbert’s projective metrics with respect to some cone of functions in settings where the marginal distributions have sufficiently light tails compared to the growth of the cost function. These results also imply the exponential convergence of the total variation distance. On normed finite dimensional spaces, the main condition in  [eckstein2025hilbert] is that $W$ grows faster that a polynomial of order $p$ and the tails of both marginals decay faster that $e^{-r^{p+q}}$ for some $q>0$. This condition doesn’t cover Gaussian models.
As underlined in [adm-24], Sinkhorn semigroups belong to the class of time varying Markov processes sharing a common invariant target measure at each individual step. These classes of time varying Markov semigroups arise in a variety of areas in applied mathematics including on nonlinear filtering, physics and molecular chemistry, see for instance [dpa, dm-04, dg-01, horton-dp, saloff-zuniga, saloff-zuniga-2, saloff-zuniga-3] and references therein.
The stability analysis of time inhomogenous Markov processes is a broad and notoriously difficult subject, much more difficult to handle than their time homogeneous counterparts; this is because the operators may drastically change during the semigroup evolution.
The articles [douc, saloff-zuniga, saloff-zuniga-2] provide several successful probabilistic and analytical tools, including coupling methods, spectral analysis as well as functional inequalities such as Nash or log-Sobolev inequalities. Sinkhorn semigroups being defined sequentially by nonlinear conjugate transformations, the application of these refined analytical techniques in this context is rather challenging. There is a need for new tools in order to handle these models. In this paper, we introduce a novel approach to Sinkhorn stability analysis that relies on Lyapunov techniques and contraction coefficients on weighted Banach spaces.
1.3 Contributions

### v3

@@ Abstract.
We develop a novel stability theory for Sinkhorn semigroups based on Lyapunov techniques and quantitative contraction coefficients, and establish exponential convergence of Sinkhorn iterations on weighted Banach spaces. This operator-theoretic framework yields explicit exponential decay rates of Sinkhorn iterates toward Schrödinger bridges with respect to a broad class of $\phi$-divergences and Kantorovich-type distances, including relative entropy, squared Hellinger integrals, $\alpha$-divergences, weighted total variation norms, and Wasserstein distances. To the best of our knowledge, these results provide the first systematic contraction inequalities of this kind for entropic transport and the Sinkhorn algorithm.
@@
We further introduce Lyapunov contraction principles under minimal regularity assumptions, leading to quantitative exponential stability estimates for a large family of Sinkhorn semigroups. The framework applies to models with polynomially growing potentials and heavy-tailed marginals on general normed spaces, as well as to more structured boundary state-space models, including semicircle transitions and Beta, Weibull, and exponential marginals, together with semi-compact settings. Finally, our approach extends naturally to statistical finite mixtures of such models, including kernel-based density estimators arising in modern generative modeling.
Key words and phrases:
@@
Entropic optimal transport, Sinkhorn semigroups, Schrödinger bridges, contraction coefficients, weighted total variation norms, Lyapunov functions, Kantorovich and entropy criteria, Wasserstein semi-distances.
2020 Mathematics Subject Classification:
Primary 37M25, 49Q22, 47H09, 60J20; secondary: 60J05, 94A17.
1. Introduction
1.1. Entropic optimal transport
@@
Let $(\mathbb{X},\lambda)$ and $(\mathbb{Y},\nu)$ be locally compact Polish spaces equipped with locally bounded positive measures $\lambda$ and $\nu$. We consider probability measures of the form
(1.1) $\lambda_{U}(dx):=e^{-U(x)}~\lambda(dx)\quad\mbox{\rm and}\quad\nu_{V}(dy):=e^{-V(y)}~\nu(dy)$
@@
where $U:\mathbb{X}\to\mathbb{R}$ and $V:\mathbb{Y}\to\mathbb{R}$ are measurable potential functions chosen so that $\lambda_{U}$ and $\nu_{V}$ are normalized. Let $\Pi_{U,V}$ be the set of probability measures $\mathcal{P}$} on $(\mathbb{X}\times\mathbb{Y})$, with prescribed marginals $(\lambda_{U},\nu_{V})$. We introduce a reference Markov kernel $\mathcal{Q}$} from $\mathbb{X}$ to $\mathbb{Y}$ of the form
(1.2) ${\mathcal{Q}}(x,dy)=e^{-W(x,y)}~\nu(dy)\quad\mbox{\rm and set}\quad{\mathcal{R}}(y,dx):=e^{-W(x,y)}~\lambda(dx).$
where $W:\mathbb{X}\times\mathbb{Y}\to\mathbb{R}$ is a measurable transition potential. The associated reference probability measure on $\mathbb{X}\times\mathbb{Y}$ is defined by
(1.3) ${\mathcal{P}}_{0}(d(x,y)):=(\lambda_{U}\times{\mathcal{Q}})(d(x,y)):=\lambda_{U}(dx)~{\mathcal{Q}}(x,dy).$
@@
The entropic optimal transport problem, also known as the Schrödinger bridge problem, with respect to the reference measure ${\mathcal{P}}_{0}$ consists in finding
@@
(1.4) $\mathbb{P}=\operatorname*{arg\,min}_{{\mathcal{P}}\,\in\,\Pi_{U,V}}\mbox{\rm Ent}({\mathcal{P}}~|~{\mathcal{P}}_{0}),\quad\mbox{\rm with the relative entropy}\quad\text{Ent}({\mathcal{P}}~|~{\mathcal{P}}_{0}):=\int\log\frac{d{\mathcal{P}}}{d{\mathcal{P}}_{0}}~d{\mathcal{P}},$
@@
where $d{\mathcal{P}}/d{\mathcal{P}}_{0}$ denotes the Radon-Nikodym derivative of $\mathcal{P}$} with respect to ${\mathcal{P}}_{0}$. We implicitly assume that there exists at least one ${\mathcal{P}}\in\Pi_{U,V}$ such that $\mbox{\rm Ent}({\mathcal{P}}|{\mathcal{P}}_{0})<\infty$. This condition ensures the existence of a solution $\mathbb{P}$ to (1.4) (cf. the seminal article by Csiszár [csiszar-2], as well as Section 6 in the Lecture Notes by Nutz [nutz], and also the survey article by Léonard [leonard] and references therein). Note that the product distribution
@@
trivially belongs to $\Pi_{U,V}$. Moreover, if ${\mathcal{P}}=(\lambda_{U}\times{\mathcal{K}})\,\in\,\Pi_{U,V}$ for some Markov kernel ${\mathcal{K}}(x,dy)$, then
$\frac{d{\mathcal{P}}}{d{\mathcal{P}}_{0}}(x,y)=\frac{d(\lambda_{U}\otimes\nu_{V})}{d{\mathcal{P}}_{0}}(x,y)\times\frac{d{\mathcal{P}}}{d(\lambda_{U}\otimes\nu_{V})}(x,y)$
which yields the decomposition
$\mbox{\rm Ent}({\mathcal{P}}~|~{\mathcal{P}}_{0})=\int~W(x,y)~{\mathcal{P}}(d(x,y))+\mbox{\rm Ent}({\mathcal{P}}~|~\lambda_{U}\otimes\nu_{V})-\int V(y)\nu_{V}(dy).$
@@
Choosing a transition potential of the form $W(x,y)={\mathcal{W}}(x,y)/t$, for some parameter $t>0$, problem (1.4) becomes equivalent to the regularized optimal transport problem
@@
(1.5) $\mathbb{P}=\operatorname*{arg\,min}_{{\mathcal{P}}\,\in\,\Pi_{U,V}}\left(\int~{\mathcal{W}}(x,y)~{\mathcal{P}}(d(x,y))+t~\mbox{\rm Ent}({\mathcal{P}}~|~\lambda_{U}\otimes\nu_{V})\right).$
@@
In the optimal transport literature, $\mathcal{W}$} is referred to as the cost function and $t>0$ as the entropic regularization parameter. The limiting case $t=0$ formally corresponds to the classical Kantorovich optimal transport problem with cost $\mathcal{W}$}. For further background and developments, we refer to [dobrushin2006perturbation, rachev1985monge, villani2008optimal] and the references therein.
1.2. Background
@@
The Schrödinger bridge problem (1.4) consists in constructing a random evolution that minimizes the relative entropy with respect to a given reference model while matching prescribed initial and final marginal distributions. This convex optimization problem can be interpreted as an entropic regularization of the Monge–Kantorovich optimal transport problem; we refer to [leonard, nutz, peyre] for recent surveys and comprehensive bibliographies.
@@
Entropic optimal transport and Schrödinger bridges have become central tools in a wide range of applications, including generative modeling and machine learning [arjovsky, doucet-bortoli, kolouri, peyre, rioux2023semi, rioux2024entropic], statistical barycenters [agueh, andersen, bigot, cramer-2, cuturi-w, loubes], economics [bojitov], computer vision [dominitz, solomon], and control theory [chen, chen-2].
Under mild integrability conditions, the Schrödinger bridge problem admits a unique solution that can be characterized as the fixed point of the Sinkhorn equations, also known as the iterative proportional fitting procedure. We refer to the pioneering works [deming, fortet, kruithof, sinkhorn-2, sinkhorn-3] and to more recent developments in [chen, cuturi, leonard-2] for historical background and modern perspectives. Sinkhorn bridges involve nonlinear integral operators acting on infinite-dimensional function spaces, and explicit solutions are generally unavailable. In finite state spaces, the Sinkhorn algorithm reduces to alternating row and column normalizations of matrices. For Gaussian models, however, Sinkhorn equations can be solved analytically using sequential Bayesian conjugacy arguments [adm-24, cramer-2, cramer-3, janadi].
@@
Sub-linear convergence rates of Sinkhorn equations have also been developed in the articles [alts-2017, chak-2018, dvu-2017]. Linear rates on non-necessarily compact spaces were first obtained by Léger in [leger] using elegant gradient descent and Bregman divergence techniques (see also the recent articles [doucet-bortoli, karimi-2024]). Refined convergence rates at least one order faster for a general class of models including sub-Gaussian target marginals has also been developed in [promit-2022] under some exponential moments estimates on Schrödinger potentials. Most existing results on the exponential convergence of Sinkhorn iterations concern finite or compact state spaces, or bounded cost functions, typically relying on Hilbert’s projective metric and related techniques [borwein, fienberg, sinkhorn-2, sinkhorn-3, soules, chen, deligiannidis, franklin, marino]. See also [greco2023non] for refined analyses on the torus and [chizat2025sharper] for recent advances in compact settings.
As emphasized in [cohen], contraction in the Hilbert metric is a very strong property, as this metric dominates all $\phi$-divergences. While Hilbert metric contraction implies exponential decay for a broad class of divergences (including total variation, relative entropy, and Wasserstein distances), it is generally difficult to obtain sharp contraction estimates for specific divergence criteria. Moreover, even on totally bounded spaces such as $\mathbb{X}=]0,1[=\mathbb{Y}$, the cost function and Schrödinger potentials may be unbounded, and the associated Markov operators may have unit Birkhoff contraction coefficient; see Sections 3.2.1, 5.3, and 5.4.
Quadratic costs on Euclidean spaces provide another important class of unbounded models. In this setting, [adm-24] develops a refined stability analysis for Gaussian Schrödinger bridges based on matrix Riccati equations. Beyond Gaussian models, only a few recent works address exponential convergence of Sinkhorn iterations for non-compact spaces and unbounded costs, notably [durmus, chiarini, eckstein2025hilbert]. The article [durmus] establishes quantitative exponential convergence for symmetric quadratic costs and marginal distributions with asymptotically positive log-concavity, using diffusion reflection couplings and convexity propagation techniques [conforti2024weak, eberle2016reflection]. These results were further refined in [chiarini], which applies to all regularization parameters and certain perturbations of log-concave models. A central tool in [chiarini] is a stability theorem controlling the entropy of Schrödinger bridges by the entropy of their marginals, under uniform semi-concavity assumptions on unknown dual Schrödinger potentials. This approach is also limited to symmetric quadratic-type costs and does not cover important classes of nonsymmetric reference transitions, such as those arising in linear Gaussian models for Ornstein–Uhlenbeck processes and denoising diffusion models. Recent work [del2025stability] partially overcomes these limitations by establishing entropic continuity results for nonsymmetric transitions and strongly convex-at-infinity potentials.
@@
An alternative approach based on generalized Hilbert metrics for weighted function spaces is developed in [eckstein2025hilbert]. Extending results in [chen, deligiannidis] to unbounded cost functions, the recent article  [eckstein2025hilbert] presents exponential decays of these new classes of Hilbert’s projective metrics with respect to some cone of functions in settings where the marginal distributions have sufficiently light tails compared to the growth of the cost function. These results also imply the exponential convergence of the total variation distance. On normed finite dimensional spaces, the main condition in  [eckstein2025hilbert] is that $W$ grows faster that a polynomial of order $p$ and the tails of both marginals decay faster that $e^{-r^{p+q}}$ for some $q>0$. This condition doesn’t cover Gaussian models.
Finally, Sinkhorn semigroups belong to the broader class of time-inhomogeneous Markov processes that share a common invariant measure at each step [adm-24]. Stability analysis for such processes is notoriously delicate, as transition operators may vary substantially over time. These classes of time varying Markov semigroups arise in a variety of areas in applied mathematics including on nonlinear filtering, physics and molecular chemistry, see for instance [dpa, dm-04, dg-01, horton-dp, saloff-zuniga-3] and references therein. While general tools exist—such as coupling methods, spectral techniques, and functional inequalities [douc, saloff-zuniga, saloff-zuniga-2]-their direct application to Sinkhorn semigroups is challenging due to the nonlinear conjugate structure of the updates. This motivates the development of new analytical tools. In this work, we introduce a Lyapunov-based, operator-theoretic framework for the quantitative stability analysis of Sinkhorn semigroups on weighted Banach spaces.
1.3. Contributions